# GenAINet: Enabling Wireless Collective Intelligence via Knowledge Transfer and Reasoning

### 26 Feb 2024 v1

@@ Abstract
GenAI and communication networks are expected to have groundbreaking synergies in 6G. Connecting generative artificial intelligence (GenAI) agents over a wireless network can potentially unleash the power of collective intelligence and pave the way for artificial general intelligence (AGI). However, current wireless networks are designed as a “data pipe” and are not suited to accommodate and leverage the power of GenAI. In this paper, we propose the GenAINet framework in which distributed GenAI agents communicate knowledge (high-level concepts or abstracts) to accomplish arbitrary tasks. We first provide a network architecture integrating GenAI capabilities to manage both network protocols and applications. Building on this, we investigate effective communication and reasoning problems by proposing a semantic-native GenAINet. Specifically, GenAI agents extract semantic concepts from multi-modal raw data, build a knowledgebase representing their semantic relations, which is retrieved by GenAI models for planning and reasoning. Under this paradigm, an agent can learn fast from other agents’ experience for making better decisions with efficient communications. Furthermore, we conduct two case studies where in wireless device query, we show that extracting and transferring knowledge can improve query accuracy with reduced communication; and in wireless power control, we show that distributed agents can improve decisions via collaborative reasoning. Finally, we address that developing a hierarchical semantic level Telecom world model is a key path towards network of collective intelligence.
@@ II GenAI Network and Agent Architectures
Wireless GenAI agents can emulate human-like decision making process, providing a path towards sophisticated and adaptive GenAI networks. On the one hand, it can bring autonomy to communication networks, such as breaking down high-level intents into low-level sub-tasks, and assigning them to different network elements for execution. On the other hand, multiple GenAI agents can collaborate by sharing knowledge to achieve common goals. To achieve this, we propose prospect architectures for GenAI networks and GenAI devices.
@@ II-A Wireless GenAI Network Architecture
In this section, we discuss several types of GenAINet architectures. In a wireless network, a GenAI agent can potentially act as an autonomous controller in the following scenarios.
Network protocols: GenAI agents can orchestrate network resources and control network functions in radio access, core, and transport domains. It can interface with network protocols to control power, resource, traffic, etc. From the network protocol perspective, the GenAINet architecture can include:
  • Hierarchical architecture: GenAI agents can be deployed in multi-level network controllers, sending policies to the data network. For instance, a commander agent operates in the RAN intelligent controller (RIC) or centralized unit (CU), while multiple executor agents operate in the network elements (NE) or the distributed unit (DU). The interactions between agents are hierarchical, where the commander plans and assigns tasks to the executors.
  • Distributed architecture: GenAI agents can be co-located with network users and control plane functions in the 5G NR and 5G core of the 3GPP architecture. The interaction between agents is self-organized, where agents exchange past observations, actions, or strategies to collaboratively accomplish the tasks.
@@
Network applications: GenAI agents can bring autonomy to network applications, such as autonomous vehicles and robots. Wireless networks allow GenAI agents to collaboratively perform various tasks, such as remote sensing, control, and planning. From the network application perspective, the GenAINet architecture can include:
  • Independent architecture: Each application is controlled by individual GenAI agents. For example, in a cellular vehicle-to-everything (C-V2X) network, the network agents control base stations, user equipments (UEs) through RIC and the traffic agents control cars, traffic lights through intelligent traffic system (ITS).
@@
Fig. 1 illustrates the considered architectures. The hierarchical and independent architecture can be built over existing 5G wireless networks where applications are over-the-top services. On the other hand, the distributed and converged architectures are more disruptive, fully implementing the vision of the GenAINet approach. However this requires: 1) convergence of network protocols and applications, where a GenAI agent brings end-to-end autonomy from sensing and transmitting, to generating information, strategies, and decisions; 2) collaboration of multiple GenAI agents, where minimal information is exchanged to plan and execute a wide range of tasks.
Figure 1: Example of GenAINet architectures with an intelligent traffic system.
@@ II-B Wireless GenAI Agent Architecture
To enable LLMs to interact with real-world scenarios, an agent has to be of a distributed structure, decision-wise and information-wise. An autonomous agent generally comprises four components: perception, action, planning, and memory. In what follows, we explain how these components can be designed to implement the proposed vision. The proposed GenAI agent is illustrated in Fig. 2. We discuss here how these modules operate in a wireless network.
Figure 2: Wireless GenAI Agent Architecture.
Perception is the part of an agent gathering useful and relevant information to an LLM. Possible information sources for perception include: 1) information sensed from the environment, e.g., channel state information (CSI), locations of user equipment (UE) and base station (BS) in a network, or on an autonomous car, the traffic density, speed and visual data of streets; 2) information received from other agents, such as messages reporting their state or response to some requested information, or strategies learned from past decisions under different states. The information can be represented by using different modalities, from raw text, images, to abstracted graphs. The perceptor should fuse and encode raw data in a common embedding space, for the LLM to perform further decision-making.
@@ III-A Semantic knowledge representation
KG is a structured representation of relations between real objects or abstracted concepts. These relations can describe connections or causalities between entities in natural language, making it easily accessible for an LLM. Two vertices would be connected by an edge if they are correlated and the weight of the edge may indicate the degree of correlation. knowledge graph (KG) can assist an LLM to perform fast information retrieval, and efficient reasoning based on the structure of entities. Moreover, with KG GenAI agents can effectively update and transfer knowledge according to the changing tasks, making communication more efficient. It can also help LLMs to understand and generate more contextually relevant and coherent responses for complex tasks requiring factual responses.
@@
TE: Topological models, e.g., hypergraphs, simplicial complexes and cell complexes, can represent the intrinsic structure of data. They can model the low-order (vertices) and high-order (simplices, cells) relations between syntactic objects or abstracted concepts. Unlike conventional neural networks learning data structure on 1D sequence or 2D grid, topological embedding (TE) can exploit more flexible and dynamic structure. Moreover, the high-order topological structure can represent more complex and causal relations, such as different word combinations by semantic features. Therefore, TE can exploit more implicit latent structures than KG. The geometric features allows for flexible structure composition according to the changes of semantic features, which saves communication and computing resources. However TE requires extending LLMs’ architecture to integrate topological information.
@@ III-B Multi-modal semantic reasoning
Figure 3: Multi-modal semantic reasoning via planning.
@@ III-C Semantic-native GenAINet
With semantic knowledge and semantic reasoning, GenAI agents can collectively learn communication protocols and decision strategies for solving different tasks, leading to an efficient GenAINet. GenAI agents can be used for various tasks on connected devices, from generating information according to the request, to producing a decision policy for taking actions to the environment. An effective communication protocol should be part of the reasoning process, guiding agents to achieve a goal, such as the accuracy of responses, the accumulated reward of actions, the fidelity of generated content, and so on. Furthermore, the agent should effectively manage the knowledge to transmit, memorize, and update its model according to the information freshness and energy cost.
GenAINet can be implemented according to a teacher-student or a distributed paradigm. The teacher-student paradigm is applicable for communication between cloud, edge, and device agents, where different sizes of LLMs are deployed. Specifically, a teacher agent with LLM trained on universal knowledge serves as a “world model”, guiding a student agent to perform specific tasks via knowledge transfer. To reduce communication cost, the teacher agent extracts knowledge in specific domains and transfer to the student, which is efficient for devices performing specific tasks (e.g., routing, traffic control). The distributed paradigm can be used for communication between mobile devices, machines, vehicles, where agents have similar capabilities and need to collaboratively complete tasks. Each agent has a memory of emergent knowledge which is learned from its past experience. During planning, the agent can retrieve or communicate knowledge with others to optimize decisions and update local knowledge. In a long term, the agents will share a common KB and make decisions locally, where communication costs can be minimized until observing new scenarios or tasks. Compared to standard communication systems, GenAINet is resource efficient and task effective. For example, in a remote query scenario the teacher-student paradigm can reduce load and latency in sending responses; in an autonomous vehicle (AV) scenario the distributed paradigm can improve driving safety and communication reliability.
@@
There are three semantic stages of information transfer between agents. The first stage is semantic compression from multi-modal sensory data. LLM can produce semantic latent on source textual data for lossless generation [12]. With multi-modal LLMs, semantic compression can be performed on images, point-clouds, to improve field-of-view or resolution with mulit-sensor data. The generated textual and visual data can be used to inform drivers, detect objects, or used by the actor to generate driving commands, which can be achieved by goal-oriented compression. The second level is semantic knowledge, where agents exchange the abstracts and concepts from memory. This can be achieved by remote retrieval, where the agent performs semantic similarity search on the request from remote agent, and send the related information. The third level is semantic reasoning, where agents find the consequence of past actions and exchange with others. This can reduce agents’ reasoning latency and improve decision reliability. Since long-term knowledge is more generalizable than short-term memory, the communication cost can be reduced.
Figure 4: Three levels of semantic native GenAINet: compression, knowledge, and reasoning.
@@ IV Case Studies of GenAINet
In this section, we show how the GenAINet can be applied to two examples namely, telecom question answering and wireless power control. Specifically, we want to show: 1) How GenAI agents can efficiently transfer knowledge to perform on-device query; 2) How GenAI agents can collaborate with reasoning to solve a wireless network problem.
@@ IV-A Knowledge transfer for wireless device query
Question answering (QnA) is a typical application of LLMs. Deploying LLMs close to the end user can largely reduce the latency and traffic burden from massive connections. Despite that light-weight LLMs are built with efficient inference techniques, they exhibit poor performance in specialized domains compared to regular LLMs. RAGs can enhance LLM’s knowledge with external database while introducing storage cost and higher latency. Envisioned from GenAINet, we propose a semantic knowledge driven on-device query. As shown in Fig. 5, GPT-3.5-turbo serves as the cloud agent, to extract from a world library the related context for the queries and compress them into latent representations to build a semantic KB. It is then sent to the LLM on devices to answer questions using RAG. The device can feedback questions to the cloud LLM when the accuracy level drops.
We have performed experiments on the TeleQnA [13], including 10k Telecom domain QnA from research and standard materials. We extract the questions’ sources and build a large vector DB in the cloud. An initial batch of sampled questions are fed into GPT-3.5 to perform retrieval from the DB, under instruction prompts to generate contexts related to the questions, and build the semantic KB for RAG on device. Table I shows that the semantic KB significantly improves the TeleQnA accuracy in all categories compared to the base Llama-7B, by reducing 27% exchanged information compared to cloud based QnA. The example shows that an LLM can extract a compressed knowledge and assist other LLMs to complete a task which is more efficient than sending raw data.
@@
QnA category GPT-3.5-turbo accuracy Semantic RAG accuracy Reduced bits exchanged
@@ IV-B Collaborative reasoning for wireless power control
Conventional power control paradigm resorts to an optimization problem by modeling the environment and finding the best power level leading to assigned performance metric (rate, energy). With the proposed paradigm, the solving is performed by exploiting LLMs’ knowledge of wireless network (acquired during pre-training), to reason over the path from observed rate to best power. This means in particular that with GenAI agents we may release training online from observed data. Fig. 6 shows our prompt construction and an example output of explanations from Llama-2-7B; the prompt clearly shows which information is given to the LLM to perform the power control task. Nothing else is provided to the LLM.
@@
The mean absolute error of data rate of all users over the target is shown in Fig. 7. We evaluated 3 scenarios: 1) standalone agents without interactions; 2) agents share the past memory of observations and actions; 3) agents share an explanation of the past decisions. It can be seen that the standalone agents find the power optimization path towards the targeted rate after several rounds. With memory sharing, agents learn slowly to understand the propagation and interference environment and adjust power in the right direction. Finally, the explanation from other agents in previous rounds can quickly help LLMs to find better power allocation scheme. The example shows that sharing reasoning outputs can effectively help LLMs to improve decision making without online training.
Figure 6: Prompt construction and example decision explanation of an agent with a memory of observations and actions.
Figure 7: Mean absolute error of data rate towards the target over different rounds, with shared memory and reasoning.
@@ V Challenges and Opportunities
LLMs are auto-regressive generative models pre-trained on natural text to predict the next token. Therefore LLMs hardly predict high-level semantics, which introduces high information redundancy and computation cost in training and inference. Besides, LLMs’ generalization capability is usually limited in domains not included in the training data. RAG and fine-tuning are commonly used to enhance LLM’s knowledge while being expensive and inflexible. Furthermore, grounding LLM’s knowledge to real world representation is challenging. Reinforcement learning is studied to optimize LLM online, which is not suitable for computational limited devices.
@@ VI Conclusion
In this paper, we introduced the GenAINet, a knowledge driven communication and reasoning network as a promising enabler for collective intelligence. The architecture of GenAINet leverages unified LLM-powered GenAI agents to optimize both network protocols and applications. Our semantic-native framework can learn structure of semantic concept or abstract from multi-modal raw data, and build semantic knoweldgebase to achieve effective communication and reasoning. Agents utilize LLMs with minimal semantic concepts retrieved locally and remotely through semantic communications to effectively plan reasoning paths and produce effective decisions for a task. We investigated a use case of transferring semantic knowledge between teacher and student agents to improve on-device query with less communication costs, as well as an example of distributed collaborative agents with reasoning to improve wireless power control. We demonstrated that our multi-agent GenAINet can unleash the power of collective intelligence, and addressed that developing a semantic-native Telecom world model is an essential path towards network collective intelligence.

### 28 Feb 2024 v2

@@ Abstract
GenAI and communication networks are expected to have groundbreaking synergies in 6G. Connecting generative artificial intelligence (GenAI) agents over a wireless network can potentially unleash the power of collective intelligence and pave the way for artificial general intelligence (AGI). However, current wireless networks are designed as a “data pipe” and are not suited to accommodate and leverage the power of GenAI. In this paper, we propose the GenAINet framework in which distributed GenAI agents communicate knowledge (high-level concepts or abstracts) to accomplish arbitrary tasks. We first provide a network architecture integrating GenAI capabilities to manage both network protocols and applications. Building on this, we investigate effective communication and reasoning problems by proposing a semantic-native GenAINet. Specifically, GenAI agents extract semantic concepts from multi-modal raw data, build a knowledgebase representing their semantic relations, which is retrieved by GenAI models for planning and reasoning. Under this paradigm, an agent can learn fast from other agents’ experience for making better decisions with efficient communications. Furthermore, we conduct two case studies where in wireless device query, we show that extracting and transferring knowledge can improve query accuracy with reduced communication; and in wireless power control, we show that distributed agents can improve decisions via collaborative reasoning. Finally, we address that developing a hierarchical semantic level Telecom world model is a key path towards a network of collective intelligence.
@@ II GenAI Network and Agent Architectures
Wireless GenAI agents can emulate human-like decision making process, providing a path towards sophisticated and adaptive GenAI networks. It can bring autonomy to network protocols and network applications. To achieve this, we propose prospect architectures for GenAI networks and GenAI agents, as illustrated in Fig. 1.
Figure 1: Proposed wireless GenAI network and agent architecture with protocol and application management.
@@ II-A Wireless GenAI Network Architecture
In a wireless network, a GenAI agent can potentially act as an autonomous controller in the following scenarios.
Network protocols: GenAI agents can orchestrate network resources and control network functions in radio access, core, and transport domains. It can interface with network protocols to control power, resource, traffic, etc. The GenAI network protocol architecture can include:
  • Hierarchical architecture: GenAI agents can be deployed in multi-level network controllers, sending policies to the data network. For instance, a commander agent operates in the RAN intelligent controller (RIC), while multiple executor agents operate in the network elements (NE). The commander deconstruct intents into subtasks and assigns to the executors.
  • Distributed architecture: GenAI agents can be co-located with distributed NEs and manageme locally the data and control plans. The interaction between agents is self-organized, where agents exchange past decisions or knowledge to collaboratively accomplish the tasks.
@@
Network applications: GenAI agents can bring autonomy to network applications, such as autonomous vehicles and robots. Wireless networks allow GenAI agents to collaboratively perform various tasks, such as remote sensing, control, and planning. The GenAI network application architecture can include:
  • Independent architecture: Each application is controlled by individual GenAI agents. For example, in a vehicular network, the network agents control base stations, user equipments (UEs) and the traffic agents control cars, traffic lights.
@@
The hierarchical and independent architecture can be built over existing 5G wireless networks where applications are over-the-top services. On the other hand, the distributed and converged architectures are more disruptive for 6G, fully implementing the protocol and application convergence with multi-agent network.
@@ II-B Wireless GenAI Agent Architecture
To enable LLMs to interact with real-world scenarios, an agent has to be of a distributed structure, decision-wise and information-wise. An autonomous agent generally comprises four components: perception, action, planning, and memory. In what follows, we explain how these components can be designed to implement the proposed vision.
Perception is the part of an agent gathering useful and relevant information to an LLM. Possible information sources for perception include: 1) information sensed from the environment, e.g., channel state information (CSI), locations of user equipment (UE) and BS in a network, or on an autonomous car, the traffic density, speed and visual data of streets; 2) information received from other agents, such as messages reporting their state or responses to some requested information, or strategies learned from past decisions under different states. The information can be represented by using different modalities, from raw text, images, to abstracted graphs. The perceptor should fuse and encode raw data in a common embedding space, for the LLM to perform further decision-making.
@@ III-A Semantic knowledge representation
KG is a structured representation of relations between real objects or abstracted concepts. These relations can describe connections and causalities between entities in natural language, making it easily accessible for an LLM. Two vertices would be connected by an edge if they are correlated and the weight of the edge may indicate the degree of correlation. knowledge graph (KG) can assist an LLM to perform fast information retrieval, and efficient reasoning based on the structure of entities. Moreover, KG can also help LLMs to understand and generate more contextually relevant and coherent responses for complex tasks requiring factual responses.
@@
TE: Topological models, e.g., hypergraphs, simplicial complexes and cell complexes, can represent the intrinsic structure of data. They can model the low-order (vertices) and high-order (simplices, cells) relations between syntactic objects or abstracted concepts. Unlike conventional neural networks learning data structure on 1D sequence or 2D grid, topological embedding (TE) can exploit more flexible and dynamic structure. Moreover, the high-order topological structure can represent more complex and causal relations, such as different word combinations by semantic features. Therefore, TE can exploit more implicit latent structures than KG. The geometric features allows for flexible structure composition according to the changes of semantic features, which saves communication and computing resources.
@@ III-B Multi-modal semantic reasoning
Figure 2: Proposed pipeline of multi-modal semantic extraction, retrieval and reasoning on GenAINet agents.
@@ III-C Semantic-native GenAINet
With semantic knowledge and semantic reasoning, GenAI agents can collectively learn communication protocols and decision strategies for solving different tasks, leading to an efficient GenAINet. GenAI agents can be used for various tasks on connected devices, from generating content to making decisions. An effective communication protocol should be part of the reasoning process, guiding agents to achieve a goal, such as the accuracy of responses, the accumulated reward of actions, and so on. Furthermore, an agent should effectively manage the knowledge to transmit, memorize, and update its model according to the information freshness and energy cost.
GenAINet can be implemented in a teacher-student or a distributed paradigm. The teacher-student paradigm is applicable for communication between cloud, edge, and device agents, where different sizes of LLMs are deployed. Specifically, a teacher agent with LLM trained on universal knowledge guides a student agent to perform specific tasks via knowledge transfer. To reduce communication cost, the teacher agent extracts knowledge in specific domains and transfer to the student, which is efficient for devices performing specific tasks (e.g., routing, traffic control). The distributed paradigm can be used for communication between mobile devices, machines, vehicles, where agents have similar capabilities and need to collaboratively complete tasks. Each agent has a memory of emergent knowledge which is learned from its past experience. During planning, the agent can retrieve or communicate knowledge with others to optimize decisions and update local knowledge. In a long term, the agents will share a common KB and make decisions locally, where communication costs can be minimized until observing new scenarios or tasks. Compared to standard communication systems, GenAINet is resource efficient and task effective. For example, in a remote query scenario the teacher-student paradigm can reduce load and latency in sending responses; in an autonomous vehicle (AV) scenario the distributed paradigm can improve driving safety and communication reliability.
@@
We shown here three typical use cases of semantic communication between car agents. The first case is semantic compression from multi-modal sensory data. LLMs can produce semantic latent on source textual data for lossless generation [12]. With multi-modal LLMs, semantic compression can be used to gather information efficiently from remote sensor with different field-of-views or resolutions, to improve downstream tasks such as accident detection, navigation. The second case is collaobrative knowledge, where agents exchange the abstracts and concepts from memory. This can be achieved by remote retrieval, where the agent performs semantic similarity search on the request from remote agent, and send the related information. The third case is semantic reasoning, where agents find the consequence of past actions and exchange with others. This can reduce agents’ reasoning latency and improve decision reliability. Since long-term knowledge is more generalized than short-term experience, the communication cost can be reduced.
Figure 3: Three use case of semantic native GenAINet: compression, control, and reasoning.
@@ IV Case Studies of GenAINet
In this section, we show how the GenAINet can be applied to two examples namely, wireless device query and wireless power control. Specifically, we want to show: 1) How GenAI agents can efficiently transfer knowledge to perform on-device query; 2) How GenAI agents can collaborate with reasoning to solve a wireless network problem.
@@ IV-A Knowledge transfer for wireless device query
Question answering (QnA) is a typical application of LLMs. Deploying LLMs close to the end user can largely reduce the latency and traffic burden from massive connections. Despite that light-weight LLMs are built with efficient inference techniques, they exhibit poor performance in specialized domains compared to regular LLMs. RAGs can enhance LLM’s knowledge with external database while introducing storage cost and higher latency. Envisioned from GenAINet, we propose a semantic knowledge driven on-device query. As shown in Fig. 4, In the cloud agent, its LLM extracts the related context for the queries from cloud data and compress them into latent representations to build a semantic KB. It is then sent to the LLM on a device to answer questions using RAG. The device can feedback questions to improve semantic extraction.
We have performed experiments on the TeleQnA [13], including 10k Telecom domain QnA from research and standard materials. We extract the questions’ sources and build a large vector DB in the cloud. An initial batch of sampled questions are fed into GPT-3.5 to perform retrieval from the DB, under instruction prompts to generate contexts related to the questions, and build the semantic KB for RAG on device. Table I shows that the semantic KB significantly improves the TeleQnA accuracy in all categories compared to the base Llama-7B, by reducing 27% exchanged information compared to cloud based QnA. The use case shows that an LLM can extract semantic knowledge and assist other LLMs to complete a task which is more efficient than sending raw data.
@@
QnA category GPT-3.5-turbo accuracy Semantic RAG accuracy Reduced bits exchange
@@ IV-B Collaborative reasoning for wireless power control
Conventional power control paradigm resorts to an optimization problem by modeling the environment and finding the best power level leading to assigned performance metric (rate, energy). With the proposed paradigm, the solving is performed by exploiting LLMs’ knowledge of wireless network (acquired during pre-training), to reason over the path from observed rate to best power. This means in particular that with GenAI agents we may release online training.
@@
The mean absolute error of data rate of all users towards the target is shown in Fig. 5. We evaluated 3 scenarios: 1) standalone agents without interactions; 2) agents share memory of past observations and actions; 3) agents share an explanation of the past decisions. It can be seen that the standalone agents find the power optimization path towards the targeted rate after several rounds. With memory sharing, agents learn slowly to understand the propagation and interference environment and adjust their power in the right direction. Finally, the explanation from other agents in previous rounds can help LLMs to find better power allocation scheme quickly. The example shows that sharing reasoning outputs can effectively help LLMs to improve decision making without online training.
Figure 5: Performances of data rate gap towards target with example prompts to LLM agents on power control.
@@ V Challenges and Opportunities
LLMs are auto-regressive generative models pre-trained on natural text to predict the next token. Therefore LLMs hardly predict high-level semantics, which introduces high information redundancy and computation cost in training and inference. Besides, LLMs’ generalization capability is usually limited in domains not included in the training data. RAG and fine-tuning are commonly used to enhance LLMs’ knowledge while being expensive and inflexible. Furthermore, grounding LLMs’ knowledge to real world representation is challenging. Reinforcement learning is studied to optimize LLM online, which is not suitable for computational limited devices.
@@ VI Conclusion
In this paper, we introduced the GenAINet, a knowledge driven communication and reasoning network as a promising enabler for collective intelligence. The architecture of GenAINet leverages unified LLM-powered GenAI agents to optimize both network protocols and applications. Our semantic-native framework can learn structure of semantic concept or abstract from multi-modal raw data, and build semantic knowledgebase to achieve effective communication and reasoning. Agents utilize LLMs with minimal semantic concepts retrieved locally and remotely through semantic communications to effectively plan reasoning paths and produce effective decisions for a task. We investigated a use case of transferring semantic knowledge between teacher and student agents to improve on-device query with less communication costs, as well as an example of distributed agents communicate with reasoning to improve wireless power control. We demonstrated that our multi-agent GenAINet can unleash the power of collective intelligence, and addressed that developing a semantic-native Telecom world model is an essential path towards network collective intelligence.